OpenStudy (xapproachesinfinity):

given the following diff equation \[(1+x^2)y"-4xy'+6y=0\] and \[y(x)=\sum_{0}^{\infty} a_n x^n\] find a formula for a_n

OpenStudy (amistre64):

define your derivatives and insert them

OpenStudy (xapproachesinfinity):

\[(1+x^2)\sum_{2}^{\infty}n(n-1)a_nx^{n-2}-4x\sum_{1}^{\infty}na_nx^n+6\sum_{0}^{\infty}a_nx^n\]

OpenStudy (xapproachesinfinity):

=0 that is

OpenStudy (amistre64):

the top infinity is assumed distribute now

OpenStudy (xapproachesinfinity):

i did evaluate y for n=o and n=1 to get all of the sum start at same starting point

OpenStudy (amistre64):

you have to distrubute, then align your exponents, then align your indexes

OpenStudy (xapproachesinfinity):

x^2 i took to inside as well as x

OpenStudy (xapproachesinfinity):

indexes are aligned with evaluating the other two so they start at n=2 no?

OpenStudy (amistre64):

(1+x^2)y" -4xy′ +6y =0 y'' +x^2y" -4xy′ +6y = 0 [a_n n(n-1) x^(n-2)]_2 +x^2[a_n n(n-1) x^(n-2)]_2 -4x[a_n n x^(n-1)]_1 +6[a_n x^n]_0 = 0 distribute your 'terms' [a_n n(n-1) x^(n-2)]_2 +[a_n n(n-1) x^(n+2-2)]_2 +[-4a_n n x^(n+1-1)]_1 +[6a_n x^n]_0 = 0 [a_n n(n-1) x^(n-2)]_2 +[a_n n(n-1) x^n]_2 +[-4a_n n x^n]_1 +[6a_n x^n]_0 = 0 we still have an exponent that is not in line, i always adjust to the 'lowest' exponent: n-2 < n [a_n n(n-1) x^(n-2)]_2 +[a_{n-2} (n-2)(n-2-1) x^(n-2)]_2+2 +[-4a_{n-2} (n-2) x^(n-2)]_1+2 +[6a_{n-2} x^(n-2)]_0+2 = 0 [a_n n(n-1) x^(n-2)]_2 +[a_{n-2} (n-2)(n-2-1) x^(n-2)]_4 +[-4a_{n-2} (n-2) x^(n-2)]_3 +[6a_{n-2} x^(n-2)]_2 = 0 now we can align our indexes up to 4

OpenStudy (xapproachesinfinity):

hmm did the first line yu did let me read the last lines to see where i missed

OpenStudy (amistre64):

srry its not in latex coding, but this amount of coding would lag my system to a halt

OpenStudy (xapproachesinfinity):

oh good i got it! that's what i needed i was doing by sub using a different index i can take from there since the sub didn't go well for i started evaluating to make the indices aligned

OpenStudy (xapproachesinfinity):

that's good enough for me thanks :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (xapproachesinfinity):

my prof was doing sub too, but he didn't get the answer lol

OpenStudy (amistre64):

my efforts .... the wolf says there are i's in it so i might have messed up along the way :) [a_n n(n-1) x^(n-2)]_2 +[a_{n-2} (n-2)(n-3) x^(n-2)]_4 +[-4a_{n-2} (n-2) x^(n-2)]_3 +[6a_{n-2} x^(n-2)]_2 = 0 a_2 2(2-1) x^(2-2) + a_3 3(3-1) x^(3-2) +[a_n n(n-1) x^(n-2)]_4 +[a_{n-2} (n-2)(n-3) x^(n-2)]_4 - 4a_{3-2} (3-2) x^(3-2) +[-4a_{n-2} (n-2) x^(n-2)]_4 + 6a_{2-2} x^(2-2) + 6a_{3-2} x^(3-2) +[6a_{n-2} x^(n-2)]_4 = 0 2a_2 + 6a_0 + 6a_3 x + 2a_1 x +[a_n n(n-1) x^(n-2)]_4 +[a_{n-2} (n-2)(n-3) x^(n-2)]_4 +[-4a_{n-2} (n-2) x^(n-2)]_4 +[6a_{n-2} x^(n-2)]_4 = 0 now that indexes and exponents are all aligned, we can combine the summations: 2a_2 + 6a_0 + 6a_3 x + 2a_1 x +[a_n n(n-1) + a_{n-2} (n-2)(n-3) -4a_{n-2} (n-2) +6a_{n-2} x^(n-2)]_4 = 0 +[a_n n(n-1) + a_{n-2}((n-2)(n-3-4)+6) x^(n-2)]_4 = 0 +[a_n n(n-1) + a_{n-2}((n-2)(n-7)+6) x^(n-2)]_4 = 0 a_n n(n-1) + a_{n-2}((n-2)(n-7)+6) = 0 a_n n(n-1) = -a_{n-2}((n-2)(n-7)+6) a_n = -a_{n-2} ((n-2)(n-7)+6)/(n(n-1)) ------------------------------------- 2a_2 + 6a_0 = 0 a_2 = -3 a_0 6a_3 + 2a_1 = 0 a_3 = -1/3 a_1 a_2 and a_3 are defined in terms of a_0 and a_1 a_4 = -a_{2} ((2)(-3)+6)/(4(3)) ; notice a_4 = 0 a_5 = -a_{3} ((3)(-2)+6)/(5(4)) ; notice a_5 = 0 a_6 = -a_{4} ((4)(-1)+6)/(6(5)) a_7 = -a_{5} ((5)(0)+6)/(7(6)) a_8 = -a_{6} ((6)(1)+6)/(8(7)) a_9 = -a_{7} ((7)(2)+6)/(9(8)) all out terms for n>=4 drop out, they are 0 y = a_0 + a_1 x + a_2 x^2 + a_3 x^3 + [0] y = a_0 + a_1 x + a_0 (-3)x^2 + a_1 (-1/3)x^3 y = a_0 + a_0 (-3)x^2 + a_1 x + a_1 (-1/3)x^3 y = a_0 (1-3x^2) + a_1 x(3-x^2)/3